# Written by AAAI Press Staff1

### v1

@@ Abstract
While research has focused on surfacing and auditing algorithmic bias to ensure equitable AI development, less is known about how NLP practitioners, those directly involved in dataset development, annotation, and deployment, perceive and navigate issues of NLP data equity. This study is among the first to center practitioners’ perspectives, linking their experiences to a multi-scalar AI governance framework and advancing participatory recommendations that bridge technical, policy, and community domains. Drawing on a 2024 questionnaire and focus group, we examine how U.S. based nlp data practitioners conceptualize fairness, contend with organizational and systemic constraints, and engage emerging governance efforts such as the U.S. AI Bill of Rights. Findings reveal persistent tensions between commercial objectives and equity commitments, alongside calls for more participatory and accountable data workflows. We critically engage debates on data diversity and “diversity-washing,” arguing that improving NLP equity requires structural governance reforms that support practitioner agency and community consent.
@@ Methods
Survey responses were analyzed descriptively in Microsoft PowerBI using frequency distributions, cross-tabulations, and percentages. Open-ended responses and focus group transcripts were coded thematically to identify patterns in bias mitigation experience, governance engagement, and organizational constraints. All study procedures were reviewed and approved by the University of Washington Institutional Review Board (IRB Protocol).
@@
Questionnaire Design The questionnaire comprised a total of 27 questions222(See the Appendices (A) in extended paper version), categorized into quantitative: multiple-choice (11), multi-select (5), rating scale (3), and qualitative: open-ended (8) types. Subsequently, the questionnaire explored a variety of themes that were mapped to the study research questions. The end of the questionnaire included questions on demographic information and a question indicating their interest in engaging in subsequent focus group studies.
Participant Demographic Details We collect and report context regarding demographics of study participants as a way to evaluate claims of data practitioner identity on the quality and performance of ML systems. Out of the 60 respondents who began the questionnaire, 13 were deemed ineligible or declined to complete the survey, resulting in 47 eligible responses for analysis. These respondents, with an average age of 29 (ranging from 19 to 40), held various professional roles in NLP data practices, including researcher, scientist, and engineer positions. The majority of them (48.94%) have an intermediate level of expertise for NLP data practices. Beginners account for 10.64%, Advanced participants represent 25.53%, and Experts make up 14.89% of the surveyed group. Demographic data collected included gender identity, race and ethnicity, languages spoken, and educational credentials 333(See the Appendices (C) in extended paper version). Specifically, the sample comprised 32 men, 13 women, one non-binary/non-conforming individual, and one participant who preferred not to disclose their gender identity. For those interested in participating in a drawing for a 50 USD Amazon gift card, email addresses were collected and aggregated into a participant interest pool. Participant demographic details are provided in (Table 1).
@@
Table 1: Survey demographic distribution of participants ($N=47$) by race/ethnicity, highlighting the diversity within the sample.
@@
In this section, survey questions are referenced using the notation “SQ” followed by the question number (e.g., SQ-10c). Focus group excerpts are labeled “FG” with a sequential number (e.g., FG-3). These labels correspond to survey or focus group questions. 444(See the Appendices (B) in extended paper version)
@@
The initial set of survey questions sought to understand the range of experiences of NLP practitioners in organizing, collecting, transcribing, and analyzing text or speech data for NLP datasets. The survey study starts with surfacing the various areas of experience among practitioners in [SQ-6a]. The analysis primarily indicated that the data work of most NLP practitioners was centered on data collection or annotation for training of the text / speech model ($76.60\%$), research on data practices or methods ($46.81\%$) and organizing and managing data annotation projects ($42.55\%$). These activities were particularly relevant to NLP technologies such as automated speech recognition (ASR), virtual assistants, smart home devices, and sentiment analysis. The full spectrum of NLP technologies that NLP practitioners reported working on. 555(See the Appendices (C) (Table 3) in extended paper version) However, we find that practitioners had the least experience mitigating bias in data training sets ($23.40\%$) and developing evaluation approaches to measure the effectiveness of training data (31.91%).
@@ Discussion
Our study examines how some NLP practitioners advance NLP data equity by reflecting on practices that produce representative, contextually accurate, and justice-oriented datasets to address technical mis-alignments and social inequities among diverse users. This study offers three key contributions. First, it is among the few to center the voice of NLP practitioners—capturing the lived realities of those embedded in the data production pipeline and their perceptions of equity, bias, and governance. Second, it connects these practitioner experiences directly to a multi-scalar governance framing, linking day-to-day decisions to institutional, national, and international policy contexts. Third, it advances participatory recommendations that bridge academic, industry, and civic domains, offering actionable pathways for enhancing practitioner agency and accountability.
@@
Developing robust strategies and metrics for measuring dataset diversity is essential. Further research emphasis could be placed on techniques for mitigating bias introduced during data collection. For instance (Cunningham et al. 2024) has proposed creating human-centered evaluation methods for better assessing AI’s sociocultural consequences in language technologies on underrepresented language communities. How are NLP professionals currently measuring the diversity of their datasets? Are there limitations to these metrics? Can new metrics be developed to provide a more comprehensive picture of dataset diversity? Simulations and user experience case studies can be used to evaluate the effectiveness of different bias mitigation strategies in various NLP data collection scenarios.
@@
Yet, in industry settings, such collaboration often depends on project leaders or even organizations who may de-prioritize public engagement due to perceived effort, costs, or misconceptions about skills necessary for data work. Investing in infrastructure to upskill community-based experts is often not considered. Although companies like Google, Meta, and Microsoft have launched expert-based external community-facing advisory boards (Smith-Loud et al. 2023), these mechanisms rarely involve casual experts whose language communities have been disproportionately impacted by NLP tools among other ML technologies.

### v2

@@ Abstract
While research has focused on surfacing and auditing algorithmic bias to ensure equitable AI development, less is known about how NLP practitioners, those directly involved in dataset development, annotation, and deployment, perceive and navigate issues of NLP data equity. This study is among the first to center practitioners’ perspectives, linking their experiences to AI governance frameworks and advancing participatory recommendations that bridge technical, policy, and community domains. Drawing on a 2024 questionnaire and focus group, we examine how U.S. based NLP data practitioners conceptualize fairness, contend with organizational and systemic constraints, and engage emerging governance efforts such as the U.S. AI Bill of Rights. Findings reveal persistent tensions between commercial objectives and equity commitments, alongside calls for more participatory and accountable data workflows. We critically engage debates on data diversity and “diversity-washing,” arguing that improving NLP equity requires structural governance reforms that support practitioner agency and community consent.
@@ Introduction
*An extended version of this paper, which includes supplementary materials, technical appendices, and expanded methodological details, is available on arXiv222navigate to https://jaylcunningham.com/work for the arXiv web-link. under a Creative Commons Attribution (CC BY) license.
@@ Methods
Survey responses were analyzed descriptively in Microsoft PowerBI using frequency distributions, cross-tabulations, and percentages. Open-ended responses and focus group transcripts were coded thematically to identify patterns in bias mitigation experience, governance engagement, and organizational constraints.
All study procedures were reviewed and approved by the University of Washington Institutional Review Board (IRB Protocol).
@@
Questionnaire Design The questionnaire comprised a total of 27 questions333(See the Appendices (A) in extended paper version), categorized into quantitative (i.e., 11 multiple-choice, 5 multi-select, 3 rating scale questions) and qualitative: (i.e., 8 open-ended questions). Subsequently, the questionnaire explored a variety of themes that were mapped to the study’s research questions. The questionnaire concluded with questions on demographic information and their interest in engaging in subsequent focus group studies.
Participant Demographic Details We collect and report context regarding demographics of study participants as a way to evaluate claims of data practitioner identity on the quality and performance of ML systems. Out of the 60 respondents who began the questionnaire, 13 were deemed ineligible or declined to complete the survey, resulting in 47 eligible responses for analysis. These respondents, with an average age of 29 (ranging from 19 to 40), held various professional roles in NLP data practices, including researcher, scientist, and engineer positions. The majority of them (48.94%) have an intermediate level of expertise for NLP data practices. Beginners account for 10.64%, Advanced participants represent 25.53%, and Experts make up 14.89% of the surveyed group. Demographic data collected included gender identity, race and ethnicity, languages spoken, and educational credentials 444(See the Appendices (C) in extended paper version). Specifically, the sample comprised 32 men, 13 women, one non-binary/non-conforming individual, and one participant who preferred not to disclose their gender identity. For those interested in participating in a drawing for a 50 USD Amazon gift card, email addresses were collected and aggregated into a participant interest pool. Participant demographic details are provided in (Table 1).
@@
Table 1: Survey participant demographic distribution ($N=47$) by race/ethnicity, highlighting the sample diversity.
@@
In this section, survey questions are referenced using the notation “SQ” followed by the question number (e.g., SQ-10c). Focus group excerpts are labeled “FG” with a sequential number (e.g., FG-3). These labels correspond to survey or focus group questions. 555(See the Appendices (B) in extended paper version)
@@
The initial set of survey questions sought to understand the range of experiences of NLP practitioners in organizing, collecting, transcribing, and analyzing text or speech data for NLP datasets. The survey study starts with surfacing the various areas of experience among practitioners in [SQ-6a]. The analysis primarily indicated that the data work of most NLP practitioners was centered on data collection or annotation for training of the text / speech model ($76.60\%$), research on data practices or methods ($46.81\%$) and organizing and managing data annotation projects ($42.55\%$). These activities were particularly relevant to NLP technologies such as automated speech recognition (ASR), virtual assistants, smart home devices, and sentiment analysis. The full spectrum of NLP technologies that NLP practitioners reported working on. 666(See the Appendices (C) (Table 3) in extended paper version) However, we find that practitioners had the least experience mitigating bias in data training sets ($23.40\%$) and developing evaluation approaches to measure the effectiveness of training data (31.91%).
@@ Discussion
Our study examines how NLP practitioners advance NLP data equity by reflecting on practices that produce representative, contextually accurate, and justice-oriented datasets to address technical mis-alignments and social inequities among diverse users. This study offers three key contributions. First, it is among the few to center the voice of NLP practitioners – capturing the lived realities of those embedded in the data production pipeline and their perceptions of equity, bias, and governance. Second, it connects these practitioner experiences directly to a multi-scalar governance framing, linking day-to-day decisions to institutional, national, and international policy contexts. Third, it advances participatory recommendations that bridge academic, industry, and civic domains, offering actionable pathways for enhancing practitioner agency and accountability.
@@
Developing robust strategies and metrics for measuring dataset diversity is essential. Further research emphasis could be placed on techniques for mitigating bias introduced during data collection. For instance, Cunningham et al. (2024) has proposed creating human-centered evaluation methods for better assessing AI’s sociocultural consequences in language technologies on underrepresented language communities. How are NLP professionals currently measuring the diversity of their datasets? Are there limitations to these metrics? Can new metrics be developed to provide a more comprehensive picture of dataset diversity? Simulations and user experience case studies can be used to evaluate the effectiveness of different bias mitigation strategies in various NLP data collection scenarios.
@@
Yet, in industry settings, such collaboration often depends on project leaders or even organizations who may de-prioritize public engagement due to perceived effort, costs, or misconceptions about skills necessary for data work. Investing in infrastructure to upskill community-based experts is often not considered. Although companies like Google, Meta, and Microsoft have launched expert-based external community-facing advisory boards (Smith-Loud et al. 2023), these mechanisms rarely involve experts whose language communities have been disproportionately impacted by NLP tools among other ML technologies.